A solution of hydrogen peroxide is 13.7% by mass. What is the molarity of the
solution? Assume that the solution has a density of 1.01 g/mL.
1
Expert's answer
2021-09-17T02:12:59-0400

Moles of "H_2O_2=\frac{13.7}{34}=0.403moles"


"Density=1.01g\/ml=1010g\/l"


Litres of solution"=\frac{100}{1010}=0.099l"


Molarity"=\frac{0.403}{0.099}=4.07M"